# Supersymmetric Black Hole Hair and $AdS_{3}\times S^{3}$

### v1

@@ Appendix A Hair deformations using six dimensional supergravity
Bena, Giusto, Shigemori, and Warner [31] wrote a linear system of equations for supersymmetric solutions to the above set-up. The metric ansatz considered in their work is of the form
@@
where $H$ and ${\cal F}$ are functions and $\beta$, $\omega$ are one-forms on the four dimensional base with metric $ds_{4}^{2}=h_{mn}dx^{m}dx^{n}$. A priori, the functions $H$ and ${\cal F}$ and the one-forms $\beta$, $\omega$ and the metric $h_{mn}$ all depend on $v,x^{m}$ but not on $u$. We consider the case when the base metric is hyper-Kähler. Moreover, we restrict to the cases where the one-form $\beta$ and the base metric are taken to be independent of the $v$ coordinate. Under these assumptions, the exterior derivative of the one-form $\beta$ becomes self-dual on the base space
@@
and where for our simplified set-up $D\Phi=\widetilde{d}\Phi-\beta\wedge\dot{\Phi}=\widetilde{d}\Phi$. $\Theta_{1}$ and $\Theta_{2}$ are two two-forms on the base space. Physically speaking, $Z_{1}$ and $Z_{2}$ are two functions on the base space. $Z_{1}$ and $Z_{2}$ are the electric potentials and $\Theta_{1}$ and $\Theta_{2}$ are the magnetic two-forms. The BPS conditions imply that the electric potentials $Z_{1}$ and $Z_{2}$ and related to the function $H$ that appears in the metric (A.8) as
@@
We can embed $J_{L}=0$ BMPV black hole in this formalism rather easily. The black holes correspond to no $v$-dependence and no magnetic fluxes $\Theta_{1}=\Theta_{2}=0$. Furthermore, for our set-up dilaton is set to its constant asymptotic value. We can adjust this value such that $\phi=0$, i.e., $Z_{1}=Z_{2}=H$. For the non-rotating black hole the one form $\omega=0$. Equation (A.17) implies $H$ is harmonic; equation (A.20) tells us that ${\cal F}$ is harmonic. In Gibbons-Hawking coordinates (2.10) the choice $H=\psi(r)$ and ${\cal F}=\psi(r)-1$, cf. (2.4), gives us the $J_{L}=0$ BMPV black hole. We can check that $2G=d(\psi^{-1}du\wedge dv)+\star_{4}DH$ matches with $F^{(3)}=r_{0}(\star_{6}\epsilon_{3}+\epsilon_{3})$, cf. (2.12), with $\epsilon_{3}=\sin\theta dx^{4}\wedge d\theta\wedge d\phi$.

### v2

@@ Appendix A Hair deformations using six dimensional supergravity
Bena, Giusto, Shigemori, and Warner [35] wrote a linear system of equations for supersymmetric solutions to the above set-up (A.1)–(A.4). The metric ansatz considered in their work is of the form
@@
where $H$ and ${\cal F}$ are functions and $\beta$, $\omega$ are one-forms on the four dimensional base with metric $ds_{4}^{2}=h_{mn}dx^{m}dx^{n}$. A priori, the functions $H$, ${\cal F}$, the one-forms $\beta$, $\omega$, and the metric $h_{mn}$ all depend on $v,x^{m}$ but not on $u$. We consider the case when the base metric is hyper-Kähler. Moreover, we restrict to the cases where the one-form $\beta$ and the base metric are taken to be independent of the $v$ coordinate. Under these assumptions, the exterior derivative of the one-form $\beta$ becomes self-dual on the base space
@@
and where for our simplified set-up $D\Phi=\widetilde{d}\Phi-\beta\wedge\dot{\Phi}=\widetilde{d}\Phi$. $\Theta_{1}$ and $\Theta_{2}$ are two two-forms on the base space. $Z_{1}$ and $Z_{2}$ are two functions on the base space. Physically speaking, $Z_{1}$ and $Z_{2}$ are the electric potentials and $\Theta_{1}$ and $\Theta_{2}$ are the magnetic two-forms. The BPS conditions imply that the electric potentials $Z_{1}$ and $Z_{2}$ and related to the function $H$ that appears in the metric (A.8) as
@@
We can embed $J_{L}=0$ BMPV black hole in this formalism rather easily. The black holes correspond to no $v$-dependence and no magnetic fluxes $\Theta_{1}=\Theta_{2}=0$. Furthermore, for our set-up, the dilaton $\phi$ is set to its constant asymptotic value. We can adjust this value such that $\phi=0$, i.e., $Z_{1}=Z_{2}=H$. For the non-rotating black hole the one form $\omega=0$. Equation (A.17) then implies $H$ is harmonic; equation (A.20) tells us that ${\cal F}$ is harmonic. In Gibbons-Hawking coordinates (2.10) the choice $H=\psi(r)$ and ${\cal F}=\psi(r)-1$, cf. (2.4), gives us the $J_{L}=0$ BMPV black hole. We can check that $2G=d(\psi^{-1}du\wedge dv)+\star_{4}DH$ matches with $F^{(3)}=r_{0}(\star_{6}\epsilon_{3}+\epsilon_{3})$, cf. (2.12), with $\epsilon_{3}=\sin\theta dx^{4}\wedge d\theta\wedge d\phi$.